# TIF-UNIMI-2025-23

### v1

@@ 5.1 Renormalization group invariance and the renormalization group equation
The simplest application of renormalization group invariance is to to a physical observable that depends on a single scale. The prototypical example is the $R$ ratio, defined as the cross-section for an electron-positron pair to produce any hadronic final state, normalized to the QED cross-section for production of a muon-antimuon pair
@@
The numerator receives contributions to any Feynman diagram with partons in the final state, see Fig. 6.
Because $R$ is a ratio of total cross-sections, it is dimensionless, and it depends only on the center-of-mass energy squared $s$ of the $e^{+}e^{-}$ collision. Hence, at energies high enough that quark masses can be neglected, by dimensional analysis, it can be only a function of the coupling with dimensionless coefficients. In a classical theory one would conclude that these coefficients must be numerical constants, but in the quantized theory all predictions depend on a renormalization scale $\mu_{r}$, hence one can form a dimensionless ratio $s/\mu^{2}_{r}$ on which the coefficients may depend, so $R=R\left(\frac{s}{\mu_{r}^{2}},\alpha_{s}(\mu_{r})\right)$. However physical predictions cannot depend on the choice of $\mu_{r}$ and thus $R$ must satisfy the renormalization group equation
@@
namely that $R$ is a function of $\alpha_{s}(s)$ with numerical coefficients.
@@
It follows that upon Mellin transformation the cross-section can be written as the ordinary product of a partonic cross-section and the parton distributions. Writing now explicitly the dependence on all kinematic variables we have
@@
and we have written the partonic cross-section $\hat{\sigma}$ in terms of a coefficient function $C$ by factoring out its leading-order expression $\hat{\sigma}_{0}(Q^{2})$, which does not depend on the strong coupling as the leading-order process is electroweak. The coefficient function $C$ is consequently dimensionless and thus can only depend on dimensionless ratios.
@@
The condition is imposed on the logarithmic derivative because we know from Sect. 3.2 that the dependence on the factorization scale is logarithmic, and we have imposed it on the logarithm of the physical cross-section because this immediately leads a condition relating the coefficient function and the PDF:
@@
The partial derivative with respect to $\mu_{F}^{2}$ in Eq. (63) is performed for fixed $\frac{\mu_{r}}{\mu_{F}}$, in order to avoid generating large logarithms of the ratio of scales.
In the last step we have defined $\gamma_{qq}$ as the logarithmic derivative of the quark PDF, and the factor of 2 is a consequence of the fact that the Mellin transform of the luminosity Eq. (55) is the product of two quark PDFs. We have further made use of the fact that $\gamma_{qq}$ cannot depend on $Q^{2}$: hence, on dimensional grounds, it can only be a function of the strong coupling, which, in turn, is evaluated at scale $\mu_{r}$ as a consequence of the renormalization process. Comparing Eq. (63) to the collinear subtracted partonic cross-section Eq. (38) we identify $\gamma_{qq}$ with the Mellin transform Eq. (45) of the splitting function.
@@ 5.2 Collinear resummation
It is useful to rewrite the solution Eq. (67) in an equivalent form by changing variable of integration from $\mu_{F}^{2}$ to $\alpha_{s}(\mu_{F}^{2})$ using Eq. (51):
@@
where we have also written the logarithm of the coefficient function, order to ease the comparison with resummed results that we will derive in the next section. Using the leading-order beta function Eq. (51) the exponentiated collinear factor becomes
@@
Further using the expression of $\alpha_{s}(Q^{2})$ in terms of $\alpha_{s}(\mu_{F}^{2})$ which we get by integrating Eq. (51), namely
$\alpha_{s}(Q^{2})=\frac{\alpha_{s}(\mu_{F}^{2})}{1+\beta_{0}\alpha_{s}(\mu_{F}^{2})\ln\frac{Q^{2}}{\mu_{F}^{2}}},$ (70)
we get
@@
Equation (71) shows explicitly that the expression Eq. (69) sums to all orders in $\alpha_{s}(\mu_{F}^{2})$ the collinear logs of $\frac{Q^{2}}{\mu_{F}^{2}}$, while Eq. (72) shows that expanding this out to first order in $\alpha_{s}(\mu_{F}^{2})$ leads back to the single-emission result Eq. (38). Note that, because one can always express $\alpha(Q^{2})$ in terms of $\alpha(\mu_{F}^{2})$ or conversely, we can equivalently view the coefficient function as a function of $\frac{Q^{2}}{\mu_{f}^{2}}$ and either $\alpha(\mu_{F}^{2})$ or $\alpha(Q^{2})$, as Eqs. (68-70) explicitly demonstrate.
@@ 6.1 The soft limit in Mellin space
The Mellin transform $\gamma_{qq}$ of the splitting function $P_{qq}$ Eq. (40) is proportional to $\frac{1}{(1-x)_{+}}$. Hence it behaves as $\ln N$ as $N\to\infty$, plus a constant, up to terms that vanish as $N\to\infty$. Each collinear log of $Q^{2}$ which is summed by the collinear exponentiation of Eq. (72) is consequently accompanied by $\ln N$. This is an infrared log, corresponding to the $x\to 1$ limit of the splitting function. However, as discussed in Sect. 3.3, the collinear log itself is
@@ 6.2 Real-virtual factorization and the soft scale
The factorization is based on the (partly conjectural) fact that real emission from the internal lines of a diagram is regular in the infrared: i.e., in comparison to the eikonal contribution Eq. (17), it does not diverge as $\frac{1}{p\cdot k}$ as $k\to 0$. In fact, we have already tacitly assumed this to be the case when stating, towards the end of Sect. 3.1, that the argument presented there does not depend on the form of the amplitude $M_{0}$. Indeed, for a generic amplitude $M_{0}$, on top of the contribution from radiation from external lines shown in Fig. 3, there are in general further real-emission contributions coming from the internal lines of $M_{0}$, that we neglected in the argument of Sect. 3.1. The suppression of radiation from internal lines then leads to factorization because soft radiation from external lines factorizes due to the eikonal argument of Sect. 2.2. This suppression is relatively easy to prove in QED: the argument, originally given in Ref. [Yennie:1961ad], is summarized in Appendix A. A general proof in QCD is rather less straightforward, though the result is likely to hold at least with processes with a colorless final state. The result is represented pictorially for the Drell-Yan process in Fig. 7.
@@
Furthermore, recalling Sect. 3 we note that the radiation contribution $J(Q^{2},x;\epsilon)$, up to non-logarithmic contributions, can only depend on the kinematic variables $Q^{2}$ and $x$ through the upper limit ${k_{t}^{2}}_{\rm max}$ of the integration over the transverse momentum of the emitted partons. This means that $J$ actually depends on a fixed combination of the kinematic variables $Q^{2}$ and $x$, namely by ${k_{t}^{2}}_{\rm max}(Q^{2},x)$. The value of ${k_{t}^{2}}_{\rm max}$ for the Drell-Yan process, already given in Eq. (41), can be determined as follows. In the general case in which in the final state we have a gauge boson with momentum $q$ and a set of partons with total momentum $k^{\prime}$, momentum conservation implies
@@
The resummation is performed in Mellin space, where, of course, the coefficient function Eq. (60) in turn factorizes into the product of a function of $Q^{2}$ and a function that depends on a fixed combination of $Q^{2}$ and $N$. Equation (77) immediately implies that with ${k_{t}^{2}}_{\rm max}=Q^{2}\frac{(1-x)^{2}}{4x}$ the Mellin space coefficient function only depends on $N$ through the soft scale
@@ 6.3 Renormalization group improving the soft scale
$\gamma^{\rm phys}(N,\alpha_{s}(Q^{2}))\equiv Q^{2}\frac{d}{dQ^{2}}\ln C\left(\frac{Q^{2}}{\mu^{2}_{F}},N,\alpha_{s}(\mu^{2}_{F}),\frac{Q^{2}}{\mu^{2}_{F}}\right).$ (86)
@@
Exploiting the factorization Eq. (80), the physical anomalous dimension can be written as a sum of two contributions, which correspond respectively to the logarithmic derivative with respect to the physical scale of the virtual and real factors, $H$ and $J$ respectively, with the latter computed in Mellin space (and the former $N$-independent). Even though the physical anomalous dimension is finite, the real and virtual contributions to it are not separately finite, and thus his decomposition can only be performed in terms of regularized quantities, which requires introducing a scale. It is the independence of the physical anomalous dimension of this scale that leads to soft resummation, just like the independence of the physical cross-section of the factorization scale led to collinear resummation in Sect. 5.2.
Indeed, recalling the definition Eq. (60) of the coefficient function, and substituting the Mellin-space version of the factorized expression of the partonic cross-section Eq. (80) in the definition Eq. (86) of the physical anomalous dimension, we get
@@
where $\gamma^{(c)}=Q^{2}\frac{d}{dQ^{2}}\ln H$ and $\gamma^{(l)}=Q^{2}\frac{d}{dQ^{2}}\ln J$, and all expressions on the right-hand side are computed in $d$ dimensions, so necessarily $\alpha_{s}$ must be $\mu$-dependent. Also, we have made use of the fact that $H$ and $J$, and correspondingly $\gamma^{(c)}$ and $\gamma^{(l)}$, respectively depend on $Q^{2}$ and $\frac{Q^{2}}{N^{2}}$ only. Because $\gamma^{(c)}$ and $\gamma^{(l)}$ are both dimensionless, they can depend on these variables only through their ratio to the regularization scale $\mu^{2}$.
@@
$\lim_{\epsilon\to 0}\gamma^{(c)}\left(\frac{Q^{2}}{\mu^{2}},\alpha_{s}(\mu^{2});\epsilon\right)$ $=g_{0}^{(c)}\left(\alpha_{s}(Q^{2})\right)-\int_{\mu^{2}}^{Q^{2}}\frac{d\lambda^{2}}{\lambda^{2}}g\left(\alpha_{s}(\lambda^{2})\right)$ (91)
$\lim_{\epsilon\to 0}\gamma^{(l)}\left(\frac{Q^{2}/N^{2}}{\mu^{2}},\alpha_{s}(\mu^{2});\epsilon\right)$ $=g_{0}^{(l)}\left(\alpha_{s}(Q^{2}/N^{2})\right)+\int_{\mu^{2}}^{Q^{2}/N^{2}}\frac{d\lambda^{2}}{\lambda^{2}}g\left(\alpha_{s}(\lambda^{2})\right).$ (92)
@@
$\gamma^{\rm phys}\left(N,\alpha_{s}(M)\right)=g^{(c)}_{0}(\alpha_{s}(Q^{2}))+g^{(l)}_{0}\left(\alpha_{s}(Q^{2}/N^{2})\right)+\int_{Q^{2}}^{Q^{2}/N^{2}}\frac{d\lambda^{2}}{\lambda^{2}}g(\alpha_{s}(\lambda^{2})).$ (93)
This is manifestly independent of the scale $\mu$, and furthermore it shows that because $\gamma^{\rm phys}$ is finite, the function $g(\alpha_{s})$ is also finite.
@@
$\gamma^{\rm phys}\left(N,\alpha_{s}(Q^{2})\right)$ $=g^{(c)}_{0}(\alpha_{s}(Q^{2}))+g^{(l)}_{0}\left(\alpha_{s}(Q^{2})\right)$
@@
This expression of the physical anomalous dimension is our final result. Renormalization group invariance and the cancellation of infrared singularities between factorized real emission contribution, that only depends on the soft scale $\Lambda^{2}_{\rm soft}=\frac{Q^{2}}{N^{2}}$, and virtual contribution, that only depends on the hard scale $Q^{2}$, leads to a prediction of the full logarithmic dependence on the Mellin variable $N$ to all orders in $\alpha_{s}(Q^{2})$. As we shall see shortly, knowledge of the first-order coefficient in the expansion of $\bar{g}$ fully determines leading log resummation, knowledge of the next coefficient together with the first correction to $\bar{g}_{0}$ the next-to-leading log result, and so on. These coefficients can be determined by comparing to fixed-order results, thereby allowing prediction of logarithmic terms to all orders based on fixed-order knowledge.

### v2

@@ 5.1 Renormalization group invariance and the renormalization group equation
The simplest application of renormalization group invariance is to a physical observable that depends on a single scale. The prototypical example is the $R$ ratio, defined as the cross-section for an electron-positron pair to produce any hadronic final state, normalized to the QED cross-section for production of a muon-antimuon pair
@@
The numerator receives contributions from any Feynman diagram with partons in the final state, see Fig. 6.
Because $R$ is a ratio of total cross-sections, it is dimensionless, and it depends only on the center-of-mass energy squared $s$ of the $e^{+}e^{-}$ collision. Hence, at energies high enough that quark masses can be neglected, by dimensional analysis, it can be only a function of the coupling with dimensionless coefficients. In a classical theory one would conclude that these coefficients must be numerical constants, but in the quantized theory all predictions depend on a renormalization scale $\mu_{r}$, hence one can form a dimensionless ratio $s/\mu^{2}_{r}$ on which the coefficients may also depend, so $R=R\left(\frac{s}{\mu_{r}^{2}},\alpha_{s}(\mu_{r})\right)$. However physical predictions cannot depend on the choice of $\mu_{r}$ and thus $R$ must satisfy the renormalization group equation
@@
namely that $R$ is a function of $\alpha_{s}(s)$ with numerical coefficients. Integrating Eq. (51) one finds that $\alpha_{s}(s)$ in terms of $\alpha_{s}(\mu_{r}^{2})$ is given by
$\alpha_{s}(s)=\frac{\alpha_{s}(\mu_{r}^{2})}{1+\beta_{0}\alpha_{s}(\mu_{r}^{2})\ln\frac{s}{\mu_{r}^{2}}},$ (53)
so with the choice $\mu^{2}_{r}=s$ we are exploiting renormalization group invariance to sum to all orders in $\alpha_{s}(\mu_{r}^{2})$ logs of the scale ratio $\frac{s}{\mu_{F}^{2}}$, which is large when $s$ is large.
@@
It follows that upon Mellin transformation the cross-section can be written as the ordinary product of the partonic cross-section and the parton distributions. Writing now explicitly the dependence on all kinematic variables we have
@@
and we have written the partonic cross-section $\hat{\sigma}$ in terms of a coefficient function $C$ by factoring out its leading-order expression $\hat{\sigma}_{0}(Q^{2})$, which does not depend on the strong coupling as the leading-order process is electroweak. The coefficient function $C$ is consequently dimensionless and thus can only depend on dimensionless ratios of dimensionful variables, that include $Q^{2}$, and the renormalization and factorization scales that were respectively introduced in order to define the renormalized theory and to treat the collinear singularities.
@@
The condition is imposed on the logarithmic derivative because we know from Sect. 3.2 that the dependence on the factorization scale is logarithmic, and we have imposed it on the logarithm of the physical cross-section because this immediately leads a condition relating the coefficient function and the luminosity (u.e. the PDFs):
@@
Furthermore, exploiting the independence of result on $\mu_{r}$, we perform the partial derivative with respect to $\mu_{F}^{2}$ in Eq. (64) at fixed $\frac{\mu_{r}}{\mu_{F}}$, because with this choice the solution of the renormalization group equation we will simultaneously sum all large scale ratios.
In the last step we have defined $\gamma_{qq}$ as the logarithmic derivative of the quark PDF, and the factor of 2 is a consequence of the fact that the Mellin transform of the luminosity, Eq. (56), is the product of two quark PDFs. We have further made use of the fact that $\gamma_{qq}$ cannot depend on $Q^{2}$: hence, on dimensional grounds, it can only be a function of the strong coupling, which, in turn, is evaluated at scale $\mu_{r}$ as a consequence of the renormalization process. Comparing Eq. (64) to the collinear subtracted partonic cross-section given by Eq. (38), we identify $\gamma_{qq}$ with the Mellin transformed expression (45) of the splitting function.
@@ 5.2 Collinear resummation
This solution manifestly displays the independence of physical observables on the factorization scale: indeed, the dependence of $C\left(\frac{Q^{2}}{\mu^{2}_{F}},N,\alpha_{s}(Q^{2}),\frac{Q^{2}}{\mu^{2}_{F}}\right)$ Eq. (68) on the scale $\mu_{F}$ is through the lower extreme of integration in the exponential, and this exactly matches the scale dependence of the luminosity, as given by Eq. (64). Note also that the further dependence on the scale through $\alpha_{s}$, which is present in Eq. (64) because the derivative is taken at fixed $\frac{\mu_{r}}{\mu_{F}}$, has been reabsorbed when choosing the renormalization scale $\mu_{r}^{2}=Q^{2}$, so that the derivative in Eq. (67) is a partial derivative, i.e. only acts on the direct dependence of $C\left(\frac{Q^{2}}{\mu^{2}_{F}},N,\alpha_{s}(Q^{2}),\frac{Q^{2}}{\mu^{2}_{F}}\right)$ on $\frac{Q^{2}}{\mu^{2}_{D}}$, for fixed $\alpha_{s}$. This generalizes the procedure that we followed when solving the renormalization group equation (50) which we can now view as the special case of the Callan-Symanzik equation (65) in which the anomalous dimension vanishes.
@@
It is useful to rewrite the solution (68) in an equivalent form by changing variable of integration from $\mu_{F}^{2}$ to $\alpha_{s}(\mu_{F}^{2})$ using Eq. (51):
@@
where we have also written the logarithm of the coefficient function, in order to ease the comparison with resummed results that we will derive in the next section. Using the leading-order beta function Eq. (51) the exponentiated collinear factor becomes
@@
Further, using Eq. (53) to express $as(Q^{2})$ in terms of $\alpha_{s}(\mu_{F}^{2})$ we get
@@
Equation (71) shows explicitly that the expression (70) sums to all orders in $\alpha_{s}(\mu_{F}^{2})$ the collinear logs of $\frac{Q^{2}}{\mu_{F}^{2}}$: we have indeed exploited renormalization group invaraince to sum all large scale ratios. On the other hand, Eq. (72) shows that expanding this out to first order in $\alpha_{s}(\mu_{F}^{2})$ leads back to the single-emission result Eq. (38). Note that, because one can always express $\alpha_{s}(Q^{2})$ in terms of $\alpha_{s}(\mu_{F}^{2})$ or conversely, we can equivalently view the coefficient function as a function of $\frac{Q^{2}}{\mu_{f}^{2}}$ and either $\alpha_{s}(\mu_{F}^{2})$ or $\alpha_{s}(Q^{2})$, as Eqs. (69-53) explicitly demonstrate.
@@ 6.1 The soft limit in Mellin space
The Mellin transform $\gamma_{qq}$ of the splitting function $P_{qq}$, as given by Eq. (40), is proportional to $\frac{1}{(1-x)_{+}}$. It follows that for large $N$ it behaves as $\ln N$ as $N\to\infty$, plus a constant, up to terms that vanish as $N\to\infty$. Each collinear log of $Q^{2}$ which is summed by the collinear exponentiation of Eq. (72) is consequently accompanied by $\ln N$. This is an infrared log, corresponding to the $x\to 1$ limit of the splitting function. However, as discussed in Sect. 3.3, the collinear log itself is
@@ 6.2 Real-virtual factorization and the soft scale
The factorization is based on the (partly conjectural) fact that real emission from the internal lines of a diagram is regular in the infrared: i.e., in comparison to the eikonal contribution (17), it does not diverge as $\frac{1}{p\cdot k}$ as $k\to 0$. In fact, we have already tacitly assumed this to be the case when stating, towards the end of Sect. 3.1, that the argument presented there does not depend on the form of the amplitude $M_{0}$. Indeed, for a generic amplitude $M_{0}$, on top of the contribution from radiation from external lines shown in Fig. 3, there are in general further real-emission contributions coming from the internal lines of $M_{0}$, that we neglected in the argument of Sect. 3.1. The suppression of radiation from internal lines then leads to factorization, because soft radiation from external lines factorizes due to the eikonal argument of Sect. 2.2. This suppression is relatively easy to prove in QED: the argument, originally given in Ref. [19], is summarized in Appendix A. A general proof in QCD is rather less straightforward, though the result is likely to hold at least with processes with a colorless final state. The result is represented pictorially for the Drell-Yan process in Fig. 7.
@@
Furthermore, recalling the argument from Sects. 3.2-3.3, we note that the radiation contribution $J(Q^{2},x;\epsilon)$, up to non-logarithmic terms, can only depend on the kinematic variables $Q^{2}$ and $x$ through the upper limit ${k_{t}^{2}}_{\rm max}$ of the integration over the transverse momentum of the emitted partons. This means that $J$ actually depends on a fixed combination of the kinematic variables $Q^{2}$ and $x$, namely by ${k_{t}^{2}}_{\rm max}(Q^{2},x)$. The value of ${k_{t}^{2}}_{\rm max}$ for the Drell-Yan process, already given in Eq. (41), can be determined as follows. In the general case in which in the final state we have a gauge boson with momentum $q$ and a set of partons with total momentum $k^{\prime}$, momentum conservation implies
@@
The resummation is performed in Mellin space, where, of course, the coefficient function, given by Eq. (61), in turn factorizes into the product of a function of $Q^{2}$ and a function that depends on a fixed combination of $Q^{2}$ and $N$. Equation (77) immediately implies that with ${k_{t}^{2}}_{\rm max}=Q^{2}\frac{(1-x)^{2}}{4x}$ the Mellin space coefficient function only depends on $N$ through the soft scale
@@ 6.3 Renormalization group improving the soft scale
$\gamma^{\rm phys}(N,\alpha_{s}(Q^{2}))\equiv Q^{2}\frac{d}{dQ^{2}}\ln C\left(\frac{Q^{2}}{\mu^{2}_{F}},N,\alpha_{s}(Q^{2}),\frac{Q^{2}}{\mu^{2}_{F}}\right).$ (86)
@@
Exploiting the factorization Eq. (80), the physical anomalous dimension can be written as a sum of two contributions, which correspond respectively to the logarithmic derivative with respect to the physical scale of the virtual and real factors, $H$ and $J$ respectively, with the latter computed in Mellin space (and the former $N$-independent). Even though the physical anomalous dimension is finite, the real and virtual contributions to it are not separately finite, and thus this decomposition can only be performed in terms of regularized quantities, which requires introducing a scale. It is the independence of the physical anomalous dimension of this scale that leads to soft resummation, just like the independence of the physical cross-section of the factorization scale led to collinear resummation in Sect. 5.2.
Indeed, recalling the definition (61) of the coefficient function, and substituting the Mellin-space version of the factorized expression of the partonic cross-section of Eq. (80) in the definition (86) of the physical anomalous dimension, we get
@@
where $\gamma^{(c)}=Q^{2}\frac{d}{dQ^{2}}\ln H$ and $\gamma^{(l)}=Q^{2}\frac{d}{dQ^{2}}\ln J$, and all expressions on the right-hand side are computed in $d$ dimensions, so necessarily $\alpha_{s}$ must be $\mu$-dependent. Also, we have made use of the fact that $H$ and correspondingly $\gamma^{(c)}$ only depends on $Q^{2}$, while $J$, and correspondingly $\gamma^{(l)}$, only depends on $\frac{Q^{2}}{N^{2}}$. Because $\gamma^{(c)}$ and $\gamma^{(l)}$ are both dimensionless, they can depend on $Q^{2}$ or respectively $Q^{2}/N$ only through their ratio to the regularization scale $\mu^{2}$. The $N$ dependence, and thus the logs, is contained in $\gamma^{(l)}$, while $\gamma^{(c)}$ is constant, i.e. NN–independent.
@@
$\gamma^{(c)}\left(\frac{Q^{2}}{\mu^{2}},\alpha_{s}(\mu^{2});\epsilon\right)$ $=g_{0}^{(c)}\left(\alpha_{s}(Q^{2});\epsilon\right)-\int_{\mu^{2}}^{Q^{2}}\frac{d\lambda^{2}}{\lambda^{2}}g\left(\alpha_{s}(\lambda^{2})\right)$ (91)
$\gamma^{(l)}\left(\frac{Q^{2}/N^{2}}{\mu^{2}},\alpha_{s}(\mu^{2});\epsilon\right)$ $=g_{0}^{(l)}\left(\alpha_{s}(Q^{2}/N^{2});\epsilon\right)+\int_{\mu^{2}}^{Q^{2}/N^{2}}\frac{d\lambda^{2}}{\lambda^{2}}g\left(\alpha_{s}(\lambda^{2})\right).$ (92)
@@
$\gamma^{\rm phys}\left(N,\alpha_{s}(Q^{2})\right)=\lim_{\epsilon\to 0}\left[g^{(c)}_{0}(\alpha_{s}(Q^{2});\epsilon)+g^{(l)}_{0}\left(\alpha_{s}(Q^{2}/N^{2});\epsilon\right)\right]+\int_{Q^{2}}^{Q^{2}/N^{2}}\frac{d\lambda^{2}}{\lambda^{2}}g(\alpha_{s}(\lambda^{2})).$ (93)
This is manifestly independent of the scale $\mu$, and furthermore it shows that because $\gamma^{\rm phys}$ is finite, both the function $g(\alpha_{s})$, and the sum of the initial conditions $g^{(c)}_{0}(\alpha_{s}(Q^{2});\epsilon)$ and $g^{(l)}_{0}\left(\alpha_{s}(Q^{2}/N^{2});\epsilon\right)$ must also be finite.
@@
$\gamma^{\rm phys}\left(N,\alpha_{s}(Q^{2})\right)$ $=\lim_{\epsilon\to 0}\left[g^{(c)}_{0}(\alpha_{s}(Q^{2});\epsilon)+g^{(l)}_{0}\left(\alpha_{s}(Q^{2});\epsilon\right)\right]$
@@
where $\bar{g}_{0}$ and $\bar{g}$ are perturbative series in $\alpha_{s}$. This expression of the physical anomalous dimension is our final result. Renormalization group invariance and the cancellation of infrared singularities between factorized real emission contribution, that only depends on the soft scale $\Lambda^{2}_{\rm soft}=\frac{Q^{2}}{N^{2}}$, and virtual contribution, that only depends on the hard scale $Q^{2}$, leads to a prediction of the full logarithmic dependence on the Mellin variable $N$ to all orders in $\alpha_{s}(Q^{2})$. As we shall see shortly, knowledge of the first-order coefficient in the expansion of $\bar{g}$ fully determines leading log resummation, knowledge of the next coefficient together with the first correction to $\bar{g}_{0}$ determines the next-to-leading log result, and so on. These coefficients can be determined by comparing to fixed-order results, thereby allowing prediction of logarithmic terms to all orders based on fixed-order knowledge.